OpenStudy (anonymous):

find the vertex line of symmetry,the max/min value of quadratic function F9x)=-2x^2+2x+6

OpenStudy (anonymous):

for vertex use \[x=-\frac{b}{2a}=-\frac{2}{2\times -2}=\frac{1}{2}\]

OpenStudy (anonymous):

so axis of symmetry is \[y=\frac{1}{2}\]to find the second coordinate of the vertex replace x by 1/2 and see what you get. i get \[\frac{13}{2}\] so that is your maximum

OpenStudy (anonymous):

so the first vertex is 2/2x-2=1/2